David can paint a wall in 1 hour and Mike can paint the same wall in 2 hours. How long would it take David and Mike to paint this wall if they work together?
Answer by Cromlix(4381)

1/1 + 1/2 = 1/x
Common denominator = 2x
2x + x = 2
3x = 2
x = 2/3 hr
x = 40 minutes.
Hope this helps.
:-)